OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? 4x2 + 20x - 24

OpenStudy (mathlegend):

I use what is known as the AC method.

OpenStudy (mathlegend):

4x^2+20x-24 is also known as.... a^2+bx+c right?

OpenStudy (mathlegend):

a= 4 b = 20 c = -24 ac = -96 b = 20 (What multiplies to -96 but equals positive 20)

OpenStudy (mathlegend):

-4*24 = -96 -4 + 24 = 20

OpenStudy (mathlegend):

So I must replace the 20x with those two digits

OpenStudy (mathlegend):

4x^2-4x+24x-24

OpenStudy (mathlegend):

Now, take the first two numbers and factor those out.... 4x(x-1) Take the last two numbers and factor those... 24(x-1) So combined we have... 4x(x-1)+24(x-1)

OpenStudy (mathlegend):

Notice inside of the parenthses it is the same. They both are (x-1) it should always be the same.

OpenStudy (mathlegend):

Now we put together the outside terms ... (4x+24)(x-1)

OpenStudy (mathlegend):

Do you understand @ermete.mayas

OpenStudy (anonymous):

kind of... omg you should have been my teacher instead -.-

OpenStudy (mathlegend):

:)

OpenStudy (anonymous):

lOl, thank you

OpenStudy (mathlegend):

No problem.